OpenStudy (anonymous):

i need help 12 less than the product of a number and -6 then simplify

OpenStudy (a_clan):

(x).(-6) - 12

OpenStudy (anonymous):

i need help, please type out the entire problem ;)

OpenStudy (anonymous):

12 less than the product of a number and negative 6 then simplify the expression

OpenStudy (anonymous):

The product of a number and -6 is: -6x 12 less than this number is: -6x-12 The simplified version would presumably involve factoring this. What's the highest common factor of -6x, and -12?

OpenStudy (anonymous):

2? or 3?

OpenStudy (anonymous):

1 and 2 factor out a -6

OpenStudy (anonymous):

Even higher than that! (or perhaps lower, as they're both negative)

OpenStudy (anonymous):

6

OpenStudy (anonymous):

6 is the correct magnitude! Notice that both of the numbers have a negative sign though, so the 'best' factor to use would probably by -6. When you factor it out, you get:\[-6x-12=-6(\frac{-6x}{-6}+\frac{-12}{-6})\]What does that equal?

OpenStudy (anonymous):

adding -6x and -12 would equal -18x? but wouldn't that be unlike terms? and if subtracting it would -6x?

OpenStudy (anonymous):

Yeah, you can't add, or subtract them I'm afraid!\[-6x-12=-6(\frac{-6x}{-6}+\frac{-12}{-6})\]Look at those fractions, and try to simplify them.

OpenStudy (anonymous):

if i simplify -6x by 6 and 12 by also 6 it leaves me with 2?

OpenStudy (anonymous):

You're right on the second count. -12 divided by -6 is 2, but we need to work harder on the first part!\[-6x-12=-6(\frac{-6x}{-6}+\frac{-12}{-6})=-6\left(\frac{-6x}{-6}+2\right)\]What is -6x divided by -6?

OpenStudy (anonymous):

1

OpenStudy (anonymous):

What happened to the x?

OpenStudy (anonymous):

1x?

OpenStudy (anonymous):

Yarp, good job! So now you have \[-6x-12=-6(\frac{-6x}{-6}+\frac{-12}{-6})=-6\left(x+2\right)\]

OpenStudy (anonymous):

thank you very much for the help!